UT 20131025 - 20131107 LUCI2 Daytime Commissioning

Observer: LUCI-team (partly remote), DThompson
Telescope Operator: various
LBTO Support: various

Plan for the period

During the period October 25 - November 8 all remaining tests not been done on October 24 but required for a successfull start of the nighttime commissioning starting November 9 must be carried out.
These are:

Tests script to take flexure data (few times 30mins)
Carried out by KPolsterer and APramskij between Oct 25 - 27

Flex-data (N3.75/ima, N1.80/210) at FIXED SLOT 6hrs
Carried out by DThompson, fixed slot tbd by Joar Brynnel

Confirmation flex-data (N3.75/ima, N1.80/210) at FIXED SLOT 6hrs
Carried out by DThompson, fixed slot tbd by Joar Brynnel

If weather is bad during science time Oct 28 - Nov 5, the following further tests can be done in closed dome (to be coordinated via DThompson):

Detector focus BB-filters
Pupil imaging & alignment
Remaining flex-data/confirmation flex-data N1.80/200/150
Detector tests (FW, linearity, persistence) by J. Panduro
Mask repro tests (fpu-turnout/fpu-cabinet)
Rough grating calibration

If weather is good, ie dome dark, the following detector tests can be done:

Readout noise, dark current, background with blind mas, blind/blind filters

Provide details of the data taken during each of the nights individually below

-- JochenHeidt - 20 Oct 2013
Topic revision: r2 - 22 Oct 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback